one sum is invested at 5% and three times as much at 8%. the total annual income is $580. how large is each sum?
Answer by mananth(16946)

5%---- x
8% --- 3x
Income = 580
5*x + 8*3x = 580 *100
percentages have been converted
29x=58000
x=2000 @ 5 %
3*2000=6000 @ 8%